APPROXIMATION AND SUPPORT THEOREM FOR A TWO SPACE-DIMENSIONAL WAVE EQUATION

ANNIE MILLET AND MARTA SANZ-SOLÉ

Abstract. We prove a characterization of the support of the law of the solution for a stochastic wave equation with two-dimensional space variable, driven by a noise white in time and correlated in space. The result is a consequence of an approximation theorem, in the convergence of probability, for equations obtained by smoothing the random noise. For some particular classes of coefficients, aproximation in the L^{p} norm, for p ≥ 1 is also proved.